if the volume of a cylinder is 462 and the height is 12 what is the radius?
Answer by checkley77(12844)

V=pir^2h
462=3.14*r^2*12
462=37.68r^2
r^2=462/37.68
r=sqrt462/37.68
r=sqrt12.26
r=3.5 ans. for the radius of the cone.
Proof:
462=3.14*3.562*12
462~462